OpenStudy (anonymous):

solutions of equation z^2-12z+36=0

OpenStudy (anonymous):

Using the quadratic equations roots formula,\[(-b \pm \sqrt{b^2-4ac})/2a\]for\[ax^2+bx+c\] \[(-(-12) \pm \sqrt{(-12)^2-4(1)(36)})/2(1)\]\[(12 \pm \sqrt{(144-144})/2\]\[(12 \pm \sqrt{(0})/2\]\[(12 \pm 0)/2\]\[12/2\]\[z=6\]

OpenStudy (anonymous):

z^2-12z+36=0 =>z^2-6z-6z+36=0 =>z(z-6)-6(z-6)=0 =>(z-6)(z-6)=0 z=6,6